OpenStudy (anonymous):

A batch of 15 DVD players contains 4 that are defective.The DVD players are selected at random,one by one,and examined. The ones that are checked are not replaced.What is the probability that there 3 defective DVD players in the first 8 DVD players examined? Help please!

OpenStudy (kropot72):

\[P(3\ defective)=\frac{4C3\times 11C5}{15C8}=\frac{4\times8\times7\times6\times7}{15\times14\times13\times12}=you\ can\ calculate\]